If the equation (n-2)x180= sum of interior angles and when we are given the interior angles use this (n-2)x180=sum (n) but when given a sum of 900 have to use the original equation? (n-2)x180=900 180n-360=900 180n=1260 n=7
Answer by Theo(13342)

yes.

the equation is that the sum of the internal angles is equal to (n - 2) * 180 where n is the number of sides in the polygon.

in a triangle, n = 3 and the sum of the internal angles is 1 * 180 = 180.

in a quadrilateral, n = 4 and the sum of the intenal angles is 2 * 180 = 360.

if the sum of the internal angles is 900, then the equation becomes:

(n - 2) * 180 = 900.

this simplifies to 180 * n - 360 = 900

add 360 to both sides of the equation to get 180 * n = 1260.

solve for n to get n = 7.

the polygon is a 7 sided figure.

i think that's called a heptagon.

here's more than you probably wanted to know about the names of polygons.

https://www.mathsisfun.com/geometry/polygons.html